1. Find the solution to the initial value problem v = 4x3 + 1 24 - 6 ( y(1) = 2. A. y= 3 – Væ4 + x - 1 B. y= 2+Vx3 + x - 2
Answer #1

you ; Yu-a (ay-6) dy = fux?+1) dx = go by = X4+X+A as y) = 2 3 4-6(2) = 1HITA 3 U-12=2+A= A=-10 ye-by=x²+x-10 a ya_by = (x²+x